# Variational Transformer Ansatz for the Density Operator of Steady States in Dissipative Quantum Many-Body Systems

### v1

@@ III.7 Evaluation of the Optimized Ansatz
$\langle\sigma^{z}\rangle=\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{z}_{i}\rangle.$ (S51)
@@
$\langle\hat{O}\rangle=\frac{\operatorname{Tr}\bigl{(}\hat{\rho}_{\bm{\theta}}\hat{O}\bigr{)}}{\operatorname{Tr}\bigl{(}\hat{\rho}_{\bm{\theta}}\bigr{)}}.$ (S52)
@@
$\langle\hat{O}\rangle=\sum_{\bm{\alpha}}P_{\bm{\theta}}(\bm{\alpha})\,\tilde{O}_{\rho_{\bm{\theta}}}(\bm{\alpha}),$ (S53)
@@
$P_{\bm{\theta}}(\bm{\alpha})=\frac{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\alpha})}{\operatorname{Tr}\bigl{(}\hat{\rho}_{\bm{\theta}}\bigr{)}},$ (S54)
@@
$\tilde{O}_{\rho_{\bm{\theta}}}(\bm{\alpha})=\sum_{\bm{\beta}}\frac{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})}{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\alpha})}\langle\bm{\beta}|\hat{O}|\bm{\alpha}\rangle.$ (S55)
@@
$\tilde{\sigma}^{z}_{\rho_{\bm{\theta}}}(\bm{\alpha})=\sum_{\bm{\beta}}\frac{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})}{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\alpha})}\langle\bm{\beta}|\sigma^{z}_{i}|\bm{\alpha}\rangle,$ (S56)
@@
$\langle\sigma^{z}_{i}\rangle=\sum_{\alpha}P_{\bm{\theta}}(\bm{\alpha})\,\tilde{\sigma}^{z}_{\rho_{\bm{\theta}}}(\bm{\alpha}).$ (S57)
@@
$\langle\sigma^{z}_{i}\rangle=\operatorname{Tr}\bigl{(}\sigma^{z}_{i}\hat{\rho}\bigr{)}=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\sigma^{z}_{i}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}},$⟩ , (S58)
@@
$\sigma^{z}_{i}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\sum_{\bm{\alpha},\bm{\beta}}\rho_{\bm{\alpha},\bm{\beta}}\,\langle\sigma^{z}_{i}|\bm{\alpha},\bm{\beta}\rangle,⟨$ (S59)
@@
$\langle\sigma^{z}_{i}|\bm{\alpha},\bm{\beta}\rangle\equiv\langle\bm{\beta}|\sigma^{z}_{i}|\bm{\alpha}\rangle.$ (S60)
@@
In summary, after optimizing the ansatz, we compute observables (e.g., $\langle\sigma^{x}\rangle$, $\langle\sigma^{y}\rangle$, and $\langle\sigma^{z}\rangle$) by sampling from $P_{\bm{\theta}}(\bm{\alpha})$ in Eq. (S54) and evaluating the corresponding local estimators as described above. We compare our results with baselines calculated using NetKet (see Sec. III.6). For small system sizes ($N<7$), we employ exact diagonalization of the full Lindblad superoperator. For larger system sizes, we use the iterative BiCGStab method, which directly solves the steady-state equation ^$\hat{\mathcal{L}}{\rho}=0$. The excellent agreement between these measurements under our optimized ansatz and benchmark solutions (obtained via exact diagonalization for small systems or iterative solvers for larger systems) confirms that our transformer-based ansatz accurately captures the steady-state properties of open quantum systems.

### v2

@@ C.7 Evaluation of the Optimized Ansatz
$\langle\sigma^{z}\rangle=\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{z}_{i}\rangle.$ (68)
@@
Figure 8: Expectation values of the averaged magnetization components $\langle\sigma_{x}\rangle$, $\langle\sigma_{y}\rangle$, and $\langle\sigma_{z}\rangle$ as functions of the transverse magnetic field $B_{y}/\gamma$ for a dissipative Heisenberg spin chain with $N=10$ sites. The system parameters are $J_{x}/\gamma=1.4$, $J_{y}/\gamma=2.0$, $J_{z}/\gamma=1.0$, $B_{x}/\gamma=-1.0$, and $B_{z}/\gamma=0.1$, with periodic boundary conditions and uniform dissipation rate $\gamma$. The exact expectation values (black lines) were obtained using the iterative BiCGStab solver via NetKet, while the red dots represent results from our transformer density operator ansatz.
@@
$\langle\hat{O}\rangle=\frac{\operatorname{Tr}\bigl{(}\hat{\rho}_{\bm{\theta}}\hat{O}\bigr{)}}{\operatorname{Tr}\bigl{(}\hat{\rho}_{\bm{\theta}}\bigr{)}}.$ (69)
@@
$\langle\hat{O}\rangle=\sum_{\bm{\alpha}}P_{\bm{\theta}}(\bm{\alpha})\,\tilde{O}_{\rho_{\bm{\theta}}}(\bm{\alpha}),$ (70)
@@
$P_{\bm{\theta}}(\bm{\alpha})=\frac{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\alpha})}{\operatorname{Tr}\bigl{(}\hat{\rho}_{\bm{\theta}}\bigr{)}},$ (71)
@@
$\tilde{O}_{\rho_{\bm{\theta}}}(\bm{\alpha})=\sum_{\bm{\beta}}\frac{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})}{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\alpha})}\langle\bm{\beta}|\hat{O}|\bm{\alpha}\rangle.$ (72)
@@
$\tilde{\sigma}^{z}_{\rho_{\bm{\theta}}}(\bm{\alpha})=\sum_{\bm{\beta}}\frac{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})}{\rho_{\bm{\theta}}(\bm{\alpha},\bm{\alpha})}\langle\bm{\beta}|\sigma^{z}_{i}|\bm{\alpha}\rangle,$ (73)
@@
$\langle\sigma^{z}_{i}\rangle=\sum_{\alpha}P_{\bm{\theta}}(\bm{\alpha})\,\tilde{\sigma}^{z}_{\rho_{\bm{\theta}}}(\bm{\alpha}).$ (74)
@@
$\langle\sigma^{z}_{i}\rangle=\operatorname{Tr}\bigl{(}\sigma^{z}_{i}\hat{\rho}\bigr{)}=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\sigma^{z}_{i}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}},$⟩ , (75)
@@
$\sigma^{z}_{i}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\sum_{\bm{\alpha},\bm{\beta}}\rho_{\bm{\alpha},\bm{\beta}}\,\langle\sigma^{z}_{i}|\bm{\alpha},\bm{\beta}\rangle,⟨$ (76)
@@
$\langle\sigma^{z}_{i}|\bm{\alpha},\bm{\beta}\rangle\equiv\langle\bm{\beta}|\sigma^{z}_{i}|\bm{\alpha}\rangle.$ (77)
@@
In summary, after optimizing the ansatz, we compute observables (e.g., $\langle\sigma^{x}\rangle$, $\langle\sigma^{y}\rangle$, and $\langle\sigma^{z}\rangle$) by sampling from $P_{\bm{\theta}}(\bm{\alpha})$ in Eq. (71) and evaluating the corresponding local estimators as described above. We compare our results with baselines calculated using NetKet (see Sec. C.6). For small system sizes ($N<7$), we employ exact diagonalization of the full Lindblad superoperator. For larger system sizes, we use the iterative BiCGStab method, which directly solves the steady-state equation ^$\hat{\mathcal{L}}{\rho}=0$. The excellent agreement between these measurements under our optimized ansatz and benchmark solutions (obtained via exact diagonalization for small systems or iterative solvers for larger systems) confirms that our transformer-based ansatz accurately captures the steady-state properties of open quantum systems.
Appendix D More numerical results
Figure 9: Expectation values of the averaged magnetization components $\langle\sigma_{x}\rangle$, $\langle\sigma_{y}\rangle$, and $\langle\sigma_{z}\rangle$ as functions of the normalized transverse field strength $g/\gamma$ for a 2D dissipative Ising model on a $2\times 2$ square lattice with periodic boundary conditions. The system parameters are $V=2\gamma$ with uniform dissipation rate $\gamma$ across all 4 sites. Our transformer density operator ansatz (red dots) accurately captures the system’s behavior obtained by NetKet.
The 1D Heisenberg Model with larger number of spins. — To further demonstrate the scalability and robustness of our transformer density operator ansatz, we present additional results for a dissipative Heisenberg spin chain with $N=10$ sites. The system is governed by the same Hamiltonian as described in the main text:
with periodic boundary conditions and uniform dissipation rate $\gamma$ across all sites. The jump operators are chosen as $L_{j}=\sigma_{j}^{-}$ for local spin lowering at each site.
For this $N=10$ system, we use the same parameters as in the main text: interaction strengths $J_{x}/\gamma=1.4$, $J_{y}/\gamma=2.0$, and $J_{z}/\gamma=1.0$, with magnetic field components $B_{x}/\gamma=-1.0$ and $B_{z}/\gamma=0.1$. We vary the transverse magnetic field $B_{y}/\gamma$ to explore different regimes of the system’s behavior. The exact benchmark results are obtained using the iterative BiCGStab solver implemented in NetKet, which provides high-accuracy solutions for systems of this size.
As shown in Fig. 8, our transformer density operator ansatz maintains good accuracy across the entire parameter range for the $N=10$ system. This result confirms the scalability and reliability of our approach across different system sizes within the one-dimensional Heisenberg model family.
Two-dimensional Ising Model on a $2\times 2$ Lattice. — To provide additional validation of our approach for 2D systems and demonstrate consistency across different lattice sizes, here we present results for a smaller $2\times 2$ square lattice dissipative Ising model.
The Hamiltonian for this system follows the same form as described in the main text:
$H=\frac{V}{4}\sum_{\langle ij\rangle}\sigma^{z}_{i}\sigma^{z}_{j}+\frac{g}{2}\sum_{i=1}^{4}\sigma^{x}_{i},$ (79)
where the sum over nearest neighbors $\langle ij\rangle$ now includes the 4 bonds in the $2\times 2$ lattice with periodic boundary conditions. The dissipative dynamics are governed by local jump operators $L_{i}=\sigma^{-}_{i}$ with uniform dissipation rate $\gamma$ across all sites, and we set the interaction strength to $V=2\gamma$.
The results presented in Fig. 9 show the expectation values of the averaged magnetization components as functions of the normalized transverse field strength $g/\gamma$. Our transformer density operator ansatz (red dots) shows excellent agreement with the exact results (black lines) across the entire parameter range.